Benzene and toluene form nearly ideal solutions. I
Answer #1

(a) Assuming the solution is ideal, Raoult's law can be applied. Vapor pressure of a component in a solution in given by Raoult's Law

Pi = xi.P*

Pi = Partial pressure of component i in the mixture, xi = mole fraction of component i and P* = Vapour Pressure of pure component i.

Total vapour pressure of the solution is the sum of partial pressure of all the components in the solution.

Ptotal = Pbenzene + Ptoulene

Ptotal = xbenzene.P*benzene + xtoulene.P*toulene

(xbenzene = 1 - 0.6 = 0.4)

Ptotal = 0.4 x 0.136 + 0.6 x 0.0422 = 0.0544 + 0.02532 = 0.0797 atm \approx0.08 atm.

(b) Mole fraction of toulene in vapour phase = Vapour pressure of component / total vapour pressure

= 0.02532 / 0.0797 = 0.3176